BexleyBrooks:

Need help with some math lol. Alistair has a cube-shaped box that has a volume of 216 cubic inches. 1. Fill in the x's 3 divided by 216=3 divided by 6 times x times x=3 divided by x to the third power = x It is written better down in the comments so some people aren't confused

BexleyBrooks:

\[3\div216\]=\[\div 6 \times x \times x=3\div x ^{3}=x\]

curriful:

this still doesnt quite make sense..

Illuminous:

@curriful wrote:
this still doesnt quite make sense..
What doesn't make sense?

Treeko:

@illuminous wrote:
@curriful wrote:
this still doesnt quite make sense..
What doesn't make sense?
Its at over the place...

Treeko:

@bexleybrooks wrote:
\[3\div216\]=\[\div 6 \times x \times x=3\div x ^{3}=x\]
I cant help it doesn't look right, Sorry